ONE-SIXTH OF A DENTIST'S AGE 8 YEARS AGO EQUALS ONE-TENTH OF HIS AGE 8 YEARS HENCE. WHAT IS HIS AGE NOW?
Answer by amoresroy(361)

ONE-SIXTH OF A DENTIST'S AGE 8 YEARS AGO EQUALS ONE-TENTH OF HIS AGE 8 YEARS HENCE. WHAT IS HIS AGE NOW?
Let x = the age of the dentist now
(x-8)/6 = (x+8)/10
Cross multiply
10(x-8) = 6(x+8)
10x-80 = 6x + 48
4x = 128
x = 32
The age of dentist now is 32 years old.